In Desmos: type in the parent function y= 3 square root of x; y= 53 square root of x + 1


List all the transformation that occurred

1
Expert's answer
2021-03-12T07:39:12-0500
"y=5\sqrt[3]{x+1}"

Parent function "f(x)=\sqrt[3]{x}."

To graph "g(x)=\sqrt[3]{x+1}," we shift the graph of "f" to the left 1 unit.

To graph "h(x)=5\sqrt[3]{x+1}," stretch the graph of "g" vertically by a factor of "5."